Human Rotational Energy. A dancer is spinning at 7
Answer #1

a) moment of inertia of head = 0.5mr^2 = 0.5*0.07*56.5*0.08^2 = 0.012656 kg m^2
moment of inertia of trunk and legs = 0.5mr^2 = 0.5*0.8*56.5*0.24^2 = 1.30176 kg m^2
moment of inertia of arms = 0.5mr^2 = 0.5*0.13*56.5*0.6^2 = 1.3221 kg m^2
Total I = 2.636516 kg m^2

b) k = 0.5Iw^2 = 0.5*2.63*72*2pi/60 = 9.934 J